# Directional testing for one-way MANOVA in divergent dimensions

### v2

@@ Abstract
In this supplementary material, Section S1 reports the additional simulation studies for the homoscedastic one-way MANOVA. Sections S1.1–S1.3 investigate the different values of the number $p$ of variables and the number $g$ of groups. The directional test performs well. The log-likelihood ratio test breaks down even when $p$ is small. The other chi-square approximations, i.e. Bartlett correction and two Skovgaard’s modifications, become worse as $p$ increasing, especially in some extreme settings in Section S1.3. The corresponding simulation results for robustness of misspecification are reported in Section S1.4. On the other hand, additional simulation studies for the heteroscedastic one-way MANOVA are showed in Section S2.
@@ 1 Introduction
Hypothesis testing for multivariate mean vectors is a very important inferential problem in many applied research fields. Likelihood-based statistics and usual asymptotic results offer a general solution to this task in parametric models. Typically, such solutions are accurate when the model dimension $p$ and the sample size $n$ match the standard asymptotic setting, where the dimension of the parameter is considered fixed as the sample size increases. However, usual asymptotic results are no longer guaranteed when $p$ is not negligible with respect to $n$ (see for instance Jiang and Yang, 2013; Sur et al., 2019; Tang and Reid, 2020; He et al., 2021). As a simple illustration, the left panel of Figure 1 shows the empirical null distribution based on 10,000 Monte Carlo samples of $p$-values obtained using the asymptotic chi-square approximation for the likelihood ratio statistic when testing the equality of $g=4$ normal mean vectors under the assumption of common unknown covariance matrix. This problem is known as homoscedastic one-way multivariate analysis of variance (MANOVA), and the simulation setup is taken from the Pottery data in the R (R Core Team, 2023) package car (Fox et al., 2022): the group sizes are $5,2,5$ and $14$, respectively, with $n=\sum_{i=1}^{4}n_{i}=26$, and the dimension of the vectors is $p=5$. It is clear from Figure 1 that the standard approximate $p$-values obtained from the likelihood ratio statistic are far from being uniform, as opposed to the directional $p$-values proposed in this paper and shown in the right panel.
@@
Settings like the former, in which the values of $p$ and $n$ are comparable, may be framed in a $(p,n)$-divergent dimensional asymptotic setting where both $p$ and $n$ are allowed to increase. Indeed, the data dimension $p$ is related to the number of parameters, e.g., for the homoscedastic one-way MANOVA case in Section 3.1, the numbers of parameters is $gp+p(p+1)/2$. Inspired by Battey and Cox (2022), we distinguish between three asymptotic regimes: moderate dimensional, high dimensional and ultra-high dimensional. Here we do not deal with the ultra-high dimensional asymptotic regime, in which $p/n$ diverges or tends to a limit greater than one. Instead, we focus on the moderate dimensional asymptotic regime, in which $p/n\to 0$, in particular with $p=O(n^{\tau})$, $\tau\in(0,1)$, and on the high dimensional asymptotic regime, in which $p/n\to\kappa\in(0,1)$. In the moderate dimensional asymptotic regime, He et al. (2021) proved that in testing the equality of normal mean vectors with identical covariance matrix the likelihood ratio test is valid if $p/n^{2/3}\to 0$, while the analogous condition for its Bartlett corrected test is $p/n^{4/5}\to 0$. To our knowledge, no similar results are available for the heteroscedastic case.
@@
Higher-order likelihood solutions based on saddlepoint approximations might generally give substantial improvements over first-order solution, especially in high dimensions (see, e.g., Tang and Reid, 2020). Among these, directional inference on a vector parameter of interest, as developed by Davison et al. (2014) and Fraser et al. (2016), has proven to be particularly accurate when testing canonical parameters in exponential families. Its accuracy descends from that of the underlying saddlepoint approximation to the conditional density of the canonical statistic of interest. Empirical results in Davison et al. (2014) and Fraser et al. (2016) showed that directional tests are extremely accurate even in settings where the dimension of the parameter of interest, although lower than the sample size, has a comparable order. The use of a saddlepoint approximation, indeed, guarantees a fairly constant relative error.
@@
The use of a directional test may be motivated by the fact that, in standard asymptotics, the directional test is first-order equivalent to the likelihood ratio test. Yet, if first-order approximations are needed for the distribution of the latter, the directional test may be more convenient, given its improved accuracy (Skovgaard, 1988; Sartori, 2017). Moreover, McCormack et al. (2019) showed that the directional test coincides with many well-known exact tests. For example, when testing a specific value for the mean of a multivariate normal distribution, the directional test coincides with the exact Hotelling’s $T^{2}$ test. Huang et al. (2022) found other instances in which the directional $p$-value is exactly uniformly distributed. Such examples are related to several prominent inferential problems in which the independence of components or the equality of covariance matrices is tested in high dimensional multivariate normal models. Finally, Di Caterina et al. (2023) extended the accurate properties of directional tests to covariance selection in high dimensional Gaussian graphical models.
@@
Concentrating on tests for the hypothesis of equality of mean vectors in $g$ independent groups, this work makes a number of contributions to the current literature. First, under the assumptions of normality and identical unknown covariance matrix, we prove that the directional $p$-value is exactly uniformly distributed provided that $n=\sum_{i=1}^{g}n_{i}\geq p+g+1$, and coincides with the Hotelling’s $T^{2}$ test when $g=2$. For the more general case with $g$ unknown group covariance matrices, known as the Behrens-Fisher problem if $g=2$, the directional test is not exact. Still, we show by means of extensive simulation studies that the directional approach overperforms standard first-order solutions as well as other higher-order modifications (Skovgaard, 2001) in moderate dimensional settings. In addition, the robustness of the available solutions to the normality assumption is empirically investigated, considering multivariate $t$, skew-normal or Laplace true generating processes. In general, all the considered approaches rely on the assumed multivariate normal model and are not expected to be reliable if that is misspecified. Yet, our numerical results show that these deviations from normality do not affect much the behaviour of the various tests and identify the directional test as the preferable solution even under model misspecification. Finally, the proposed directional test is applied for the analysis of two datasets: one from a study on blood characteristics of male athletes, to test whether the blood characteristics of male athletes differ between endurance-related and power-related events; and one from a microarray analysis on breast cancer patients to test whether the gene expressions differ according to the grade of breast tumor.
@@
The rest of the paper is organized as follows. Section 2 presents some background information. In particular, Section 2.1 reviews some important likelihood-based statistics in exponential families, Section 2.2 reports the steps to compute the directional $p$-value, and Section 2.3 details the necessary quantities for the multivariate normal model. The main results in Section 3 are for hypotheses concerning: (i) the equality of $g$ normal mean vectors with identical covariance matrix (Section 3.1); (ii) the equality of $g$ normal mean vectors with different covariance matrices (Section 3.2). For hypothesis (i), we prove the exact uniform distribution of the directional $p$-value under the null. In Section 4 we report empirical results, both under the assumed model and under model misspecification. Two applications of the considered methodology to real data are illustrated in Section 5, while Section 6 concludes with a discussion. Some auxiliary computational results and proofs are available in Appendices Appendix A and Appendix B, respectively. Additional simulation studies can be found in the Supplementary Material.
@@ 2.1 Notation and setup
Suppose that data $y=(y_{1},\dots,y_{n})^{\top}$ are generated from the model $f(y;\theta)$ with parameter $\theta=(\theta_{1},\dots,\theta_{q})^{\top}$. The log-likelihood function is $\ell(\theta)=\ell(\theta;y)=\log f(y;\theta)$, and we are interested in testing the null hypothesis $H_{\psi}:\psi(\theta)=\psi$ on the $d$-dimensional parameter of interest $\psi$. It will often be the case that $\psi$ is a component or a reparameterization of $\theta$. Assume the partition $\theta=(\psi^{\top},\lambda^{\top})^{\top}$ holds, with $\lambda$ a $(q-d)$-dimensional nuisance parameter. Let $\hat{\theta}$ denote the maximum likelihood estimate of $\theta$ and $\hat{\theta}_{\psi}$ its constrained maximum likelihood estimate under $H_{\psi}$, i.e. $\hat{\theta}_{\psi}=(\psi^{\top},\hat{\lambda}_{\psi}^{\top})^{\top}$.
@@
when $q$ is fixed and $n\to\infty$, the statistic $W$ has a $\chi^{2}_{d}$ asymptotic null distribution with relative error of order $O(n^{-1})$. A correction of $W$ proposed by Bartlett (1937) rescales the likelihood ratio statistic by its expectation under $H_{\psi}$, that is
@@
and follows asymptotically a $\chi^{2}_{d}$ null distribution with relative error of order $O(n^{-2})$ (McCullagh, 2018, Section 7.4). More details on the expectation $E(W)$ can be found in Huang et al. (2022, Section 2.1).
Skovgaard (2001) introduced two improvements on $W$, namely
@@
which also have approximate $\chi^{2}_{d}$ distributions when the null hypothesis holds. The correction factor $\gamma$ can be found in Skovgaard (2001, eq.(13)), and is also reported in A.1 with the notation used here.
The test statistics, presented so far are omnibus measures of departure of the data from $H_{\psi}$: their $p$-value results from averaging the deviations from $H_{\psi}$ in all the potential directions. We now introduce the directional test developed by Davison et al. (2014) and Fraser et al. (2016), which measures the departure from $H_{\psi}$ along the direction indicated by the observed data.
@@ 2.2 Directional testing
maximum likelihood estimate $\hat{\varphi}=\varphi(\hat{\theta})$ and constrained maximum likelihood estimate $\hat{\varphi}_{\psi}=\varphi(\hat{\theta}_{\psi})$. Henceforth, we shall use the 0 supperscript to indicate quantities evaluated at the observed data point $y^{0}$. For computing the directional $p$-value, it is convenient to define a centered sufficient statistic at $y^{0}$, $s=u-u^{0}$, with $u^{0}=u(y^{0})$. Then, the tilted log-likelihood function of model (4) takes the form
@@
where $\ell^{0}(\theta)=\ell(\theta;y^{0})$ is the observed log-likelihood function. The saddlepoint approximation (see e.g., Pace and Salvan, 1997, Section 10) to the exponential model on $\mathbb{R}^{q}$ is
@@
The hypothesis $H_{\psi}$ specifies a value for the parameter $\psi=\psi(\varphi)$. Following Fraser et al. (2016), the reduced model in $\mathbb{R}^{d}$ is given by
@@
The directional test is constructed by defining a line $L_{\psi}^{*}$ through the observed value $s^{0}=0_{q}$ and the expected value of $s$ under $H_{\psi}$ which depends on the observed data point $y^{0}$, i.e. $s_{\psi}=-\ell_{\varphi}^{0}\left(\hat{\varphi}^{0}_{\psi}\right)=-\partial\ell^{0}\left\{\theta(\varphi)\right\}/\partial\varphi\left|{}_{\varphi=\hat{\varphi}^{0}_{\psi}}\right.$. We parameterize this line by $t\in\mathbb{R}$, namely $s(t)=s_{\psi}+t(s^{0}-s_{\psi})$. In particular, $s(0)=s_{\psi}$, corresponding to $H_{\psi}$, and $s(1)=s^{0}$, corresponding to the observed data. Then, the directional $p$-value measuring the departure from $H_{\psi}$ along the line $L_{\psi}^{*}$ is defined as the probability that $s(t)$ is as far or farther from $s_{\psi}$ than is the observed value $s^{0}$:
@@
where the denominator is a normalizing constant. The upper limit $t_{\text{sup}}$ of the integrals in (7) is the largest value of $t$ for which the maximum likelihood estimate $\hat{\varphi}(t)$, corresponding to $s(t)$, exists. See Fraser et al. (2016, Section 3) for more details.
In the particular case where $\psi$ and $\lambda$ are linear functions of the canonical parameter of an exponential family the quantity (6) does not depend on $s$ and can therefore be ignored in computing the directional $p$-value (7). Moreover, the expected value $s_{\psi}$ simplifies to $s_{\psi}=\left\{-\ell_{\psi}^{0}\left(\hat{\varphi}^{0}_{\psi}\right)^{\top},{{0}_{q-d}^{\top}}\right\}^{\top}$.
If the original model $f(y;\theta)$ is not in the exponential family, then a tangent exponential model is used instead of (4). The construction of the tangent exponential model and its saddlepoint approximation are described in Fraser et al. (2016, Appendix) (see also Davison and Reid, 2022). Since the working model in our paper is the normal model which belongs to the exponential family of distributions, this step is not needed, although saddlepoint approximation (5) will be needed when $\psi$ is not a linear function of the canonical parameter, as in the heteroscedastic on-way MANOVA of Section 3.2.
@@ 2.3 Independent groups from multivariate normal distributions
Consider $g$ independent groups, and denote by $y_{ij}$ the independent observations from the $i$th group with multivariate normal distribution $N_{p}(\mu_{i},\Lambda_{i}^{-1})$ $(i=1,\dots,g;j=1,\dots,n_{i})$. The mean vectors $\mu_{i}$ and the concentration matrices $\Lambda_{i}$, symmetric and positive definite, are assumed unknown. Let $\text{tr}(A)$ indicate the trace of a square matrix $A$, and $\text{vec}(A)$ be the vector stacking the columns of $A$ one by one. We also define the vector $\text{vech}(A)$, obtained from $\text{vec}(A)$ by eliminating all upper triangular elements of $A$ when this is symmetric. These two vectors satisfy the relationship $D_{p}\text{vech}(A)=\text{vec}(A)$, where $D_{p}$ is the so-called duplication matrix (Magnus and Neudecker, 1999, Section 3.8).
@@
where $\bar{y}_{i}=y_{i}^{\top}1_{n_{i}}/n_{i}$ with $1_{n_{i}}$ a $n_{i}$-dimensional vector of ones, $i=1,\dots,g$. In this exponential family model, the canonical parameter is $\varphi=\{\xi_{1}^{\top},\dots,\xi_{g}^{\top},$ ${\text{vech}}(\Lambda_{1})^{\top},\dots,{\text{vech}}(\Lambda_{g})^{\top}\}^{\top}=\{\mu_{1}^{\top}\Lambda_{1},\dots,\mu_{g}^{\top}\Lambda_{g},$ ${\text{vech}}(\Lambda_{1})^{\top},\dots,{\text{vech}}(\Lambda_{g})^{\top}\}^{\top}$. The log-likelihood as a function of $\varphi$ is
$\ell(\varphi)$ $=\sum_{i=1}^{g}n_{i}\xi_{i}^{\top}\bar{y}_{i}-\frac{1}{2}{\text{tr}}(\Lambda_{i}y_{i}^{\top}y_{i})+\frac{n_{i}}{2}\log|\Lambda_{i}|-\frac{n_{i}}{2}\xi_{i}^{\top}\Lambda_{i}^{-1}\xi_{i}$
$=\sum_{i=1}^{g}\xi_{i}^{\top}n_{i}\bar{y}_{i}-\text{vech}(\Lambda_{i})^{\top}\left\{\frac{1}{2}D_{p}^{\top}D_{p}\text{vech}(y_{i}^{\top}y_{i})\right\}+\frac{n_{i}}{2}\log|\Lambda_{i}|-\frac{n_{i}}{2}\xi_{i}^{\top}\Lambda_{i}^{-1}\xi_{i}.$
@@
The score function is $\ell_{\varphi}(\varphi)=\frac{\partial}{\partial\varphi}\ell(\varphi)=\{\ell_{\varphi_{1}}(\varphi_{1}),\dots,\ell_{\varphi_{g}}(\varphi_{g})\}^{\top}$, with
$\ell_{\varphi_{i}}(\varphi_{i})$ =$=\left\{\ell_{\xi_{i}}(\varphi_{i})^{\top},\ell_{\text{vech}(\Lambda_{i})}(\varphi_{i})^{\top}\right\}^{\top}$
=$=\left\{n_{i}\bar{y}_{i}^{\top}-n_{i}\xi_{i}^{\top}\Lambda_{i},\;\frac{n_{i}}{2}\text{vech}(\Lambda_{i}^{-1}-y_{i}^{\top}y_{i}/n_{i}+\Lambda_{i}^{-1}\xi_{i}\xi_{i}^{\top}\Lambda_{i}^{-1})^{\top}\right\}^{\top}.$
@@
The maximum likelihood estimates for $\mu_{i}$ and $\Lambda_{i}^{-1}$ are $\hat{\mu}_{i}=\bar{y}_{i}$ and $\hat{\Lambda}_{i}^{-1}=y_{i}^{\top}y_{i}/n_{i}-\bar{y}_{i}\bar{y}_{i}^{\top}$, respectively; hence, $\hat{\xi}_{i}=\hat{\Lambda}_{i}\hat{\mu}_{i}$, $i=1,\dots,g$. Moreover, the observed information matrix $J_{\varphi\varphi}$ can be written in a block-diagonal form, using the diagonal matrix $J_{\varphi_{i}\varphi_{i}}$ $(i=1,\dots,g)$ which can be found in Huang et al. (2022, Section 2.3). Then, the determinant of $J_{\varphi\varphi}$ is such that $|J_{\varphi\varphi}(\varphi)|\propto\prod_{i=1}^{g}|{\Lambda}_{i}^{-1}|^{p+2}$.

### v3

@@ Abstract
Section S1 provides auxiliary computational results for the correction factor $\gamma$ in skovgaard:2001’s modifications, as well as the observed information matrix operator used for testing hypothesis (8) of the main text. Section S2 presents the proofs for Lemma 1, Theorem 1 and Proposition 1. Section S3 reports the extensive simulation studies for the homoscedastic one-way MANOVA; Specifically, Sections S3.1–S3.4 examine varing dimenions $p$ and the group numbers $g$. While the directional test maintains robust performance, the log-likelihood ratio test fails even at small $p$. Other chi-square approximations, i.e. Bartlett correction and two skovgaard:2001’s modifications, deteriorate as $p$ increases, particularly in the extreme settings of Section S3.4. Sections S3.5–S3.6 assess robustness under model misspecification for high and moderate dimensional setups. Furthermore, Section S4 reports extensive simulations for the heteroscedastic one-way MANOVA. Finally, Section S5 provides results from an additional application to breast cancer microarray data.
@@ 1 Introduction
Hypothesis testing for multivariate mean vectors is a very important inferential problem in many applied research fields. Likelihood-based statistics and usual asymptotic results offer a general solution to this task in parametric models. Typically, such solutions are accurate when the model dimension $p$ and the sample size $n$ match the standard asymptotic setting, where the dimension of the parameter is considered fixed as the sample size increases. However, usual asymptotic results are no longer guaranteed when $p$ is not negligible with respect to $n$ (see for instance jiang:2013; sur2019likelihood; tang2020modified; He2020B). As a simple illustration, the left panel of Figure 1 shows the empirical null distribution based on 10,000 Monte Carlo samples of $p$-values obtained using the asymptotic chi-square approximation for the likelihood ratio statistic when testing the equality of $g=4$ normal mean vectors under the assumption of common unknown covariance matrix. This problem is known as homoscedastic one-way multivariate analysis of variance (MANOVA), and the simulation setup is taken from the Pottery data in the R (Rcode) package car (car): the group sizes are $5,2,5$ and $14$, respectively, with $n=\sum_{i=1}^{4}n_{i}=26$, and the dimension of the vectors is $p=5$. It is clear from Figure 1 that the standard approximate $p$-values obtained from the likelihood ratio statistic are far from being uniform, as opposed to the directional $p$-values proposed in this paper and shown in the right panel.
@@
Settings like the former, in which the values of $p$ and $n$ are comparable, may be framed in a $(p,n)$-divergent dimensional asymptotic setting where both $p$ and $n$ are allowed to increase. Indeed, the data dimension $p$ is related to the number of parameters, e.g., for the homoscedastic one-way MANOVA case in Section 3.1, the numbers of parameters is $gp+p(p+1)/2$. Inspired by battey2022some, we distinguish between three asymptotic regimes: moderate dimensional, high dimensional and ultra-high dimensional. Here we do not deal with the ultra-high dimensional asymptotic regime, in which $p/n$ diverges or tends to a limit greater than one. Instead, we focus on the moderate dimensional asymptotic regime, in which $p/n\to 0$, in particular with $p=O(n^{\tau})$, $\tau\in(0,1)$, and on the high dimensional asymptotic regime, in which $p/n\to\kappa\in(0,1)$. In the moderate dimensional asymptotic regime, He2020B proved that in testing the equality of normal mean vectors with identical covariance matrix the likelihood ratio test is valid if $p/n^{2/3}\to 0$, while the analogous condition for its Bartlett corrected test is $p/n^{4/5}\to 0$. To our knowledge, no similar results are available for the heteroscedastic case.
@@
Higher-order likelihood solutions based on saddlepoint approximations might generally give substantial improvements over first-order solution, especially in high dimensions (see, e.g., tang2020modified). Among these, directional inference on a vector parameter of interest, as developed by sartori:2014 and sartori:2016, has proven to be particularly accurate when testing canonical parameters in exponential families. Its accuracy descends from that of the underlying saddlepoint approximation to the conditional density of the canonical statistic of interest. Empirical results in sartori:2014 and sartori:2016 showed that directional tests are extremely accurate even in settings where the dimension of the parameter of interest, although lower than the sample size, has a comparable order. The use of a saddlepoint approximation, indeed, guarantees a fairly constant relative error.
@@
The use of a directional test may be motivated by the fact that, in standard asymptotics, the directional test is first-order equivalent to the likelihood ratio test. Yet, if first-order approximations are needed for the distribution of the latter, the directional test may be more convenient, given its improved accuracy (skogaard:1988; sartori2017introduction). Moreover, sartori:2019Ftest showed that the directional test coincides with many well-known exact tests. For example, when testing a specific value for the mean of a multivariate normal distribution, the directional test coincides with the exact Hotelling’s $T^{2}$ test. huang2021directional found other instances in which the directional $p$-value is exactly uniformly distributed. Such examples are related to several prominent inferential problems in which the independence of components or the equality of covariance matrices is tested in high dimensional multivariate normal models. Finally, claudia2023ss showed the accurate properties of directional tests for covariance selection in high dimensional Gaussian graphical models.
@@
Concentrating on tests for the hypothesis of equality of mean vectors in $g$ independent groups, this work makes a number of contributions to the current literature. First, under the assumptions of normality and identical unknown covariance matrix, we prove that the directional $p$-value is exactly uniformly distributed provided that $n=\sum_{i=1}^{g}n_{i}\geq p+g+1$, and coincides with the Hotelling’s $T^{2}$ test when $g=2$. For the more general case with $g$ unknown group covariance matrices, known as the Behrens-Fisher problem if $g=2$, the directional test is not exact. Still, we show by means of extensive simulation studies that the directional approach overperforms standard first-order solutions as well as other higher-order modifications (skovgaard:2001) in moderate dimensional settings. In addition, the robustness of the available solutions to the normality assumption is empirically investigated, considering multivariate $t$, skew-normal or Laplace true generating processes in the Supplementary Material. Finally, the proposed directional test is applied for the analysis of two datasets: one from a study on the blood characteristics of male athletes, and another, presented in the Supplementary Material, from a microarray analysis on breast cancer patients.
@@
The rest of the paper is organized as follows. Section 2 presents some background information. In particular, Section 2.1 reviews some important likelihood-based statistics in exponential families, Section 2.2 reports the steps to compute the directional $p$-value, and Section 2.3 details the necessary quantities for the multivariate normal model. The main results in Section 3 are for hypotheses concerning: (i) the equality of $g$ normal mean vectors with identical covariance matrix (Section 3.1); (ii) the equality of $g$ normal mean vectors with different covariance matrices (Section 3.2). For hypothesis (i), we prove the exact uniform distribution of the directional $p$-value under the null. In Section 4 we report empirical results under the assumed model. An application of the considered methodology to real data, specifically blood characteristics of male athletes, is illustrated in Section 5, while Section 6 concludes with a discussion. The Supplementary Material contains auxiliary computational results, proofs of the theorems, additional simulation studies, including larger group sizes and robustness analyses, as well as the second application involving breast cancer patient microarray data.
@@ 2.1 Notation and setup
Suppose that data $y=(y_{1},\dots,y_{n})^{\top}$ are generated from the model $f(y;\theta)$ with parameter $\theta=(\theta_{1},\dots,\theta_{q})^{\top}$. The log-likelihood function is $\ell(\theta)=\ell(\theta;y)=\log f(y;\theta)$, and we are interested in testing the null hypothesis $H_{\psi}:\psi(\theta)=\psi$ on the $d$-dimensional parameter of interest $\psi$. It will often be the case that $\psi$ is a component of $\theta$, possibly after a reparameterization. Assume the partition $\theta=(\psi^{\top},\lambda^{\top})^{\top}$ holds, with $\lambda$ a (qd)(q-d)-dimensional nuisance parameter. Let $\hat{\theta}$ denote the maximum likelihood estimate of $\theta$ and $\hat{\theta}_{\psi}$ its constrained maximum likelihood estimate under $H_{\psi}$, i.e., $\hat{\theta}_{\psi}=(\psi^{\top},\hat{\lambda}_{\psi}^{\top})^{\top}$.
@@
When $q$ is fixed and $n\to\infty$, the statistic $W$ has a $\chi^{2}_{d}$ asymptotic null distribution with relative error of order $O(n^{-1})$. A correction of $W$ proposed by bartlett:1937 rescales the likelihood ratio statistic by its expectation under $H_{\psi}$, that is
@@
and follows asymptotically a $\chi^{2}_{d}$ null distribution with relative error of order $O(n^{-2})$ (mccullage:2018, Section 7.4). More details on the expectation $E(W)$ can be found in huang2021directional.
skovgaard:2001 introduced two improvements on $W$, namely
@@
which also have approximate $\chi^{2}_{d}$ distributions when the null hypothesis holds. The correction factor $\gamma$ can be found in skovgaard:2001, and is also reported in S1.1 with the notation used here.
The test statistics presented so far are omnibus measures of departure of the data from $H_{\psi}$: their $p$-value results from averaging the deviations from $H_{\psi}$ in all the potential directions of the parameter space. We will now introduce the directional test developed by sartori:2014 and sartori:2016, which measures the departure from $H_{\psi}$ along the direction indicated by the observed data.
@@ 2.2 Directional testing
[…]
@@
where $\ell^{0}(\theta)=\ell(\theta;y^{0})$ is the observed log-likelihood function. The saddlepoint approximation (see, e.g., salvan:1997, Section 10) to the exponential model on $\mathbb{R}^{q}$ is
@@
The hypothesis $H_{\psi}$ specifies a value for the parameter $\psi=\psi(\varphi)$. Following sartori:2016, the reduced model in $\mathbb{R}^{d}$ is given by
@@
The directional test is constructed by defining a line $L_{\psi}^{*}$ through the observed value $s^{0}=0_{q}$ and the expected value of $s$ under $H_{\psi}$ which depends on the observed data point $y^{0}$, i.e. $s_{\psi}=-\ell_{\varphi}^{0}\left(\hat{\varphi}^{0}_{\psi}\right)=-\partial\ell^{0}\left\{\theta(\varphi)\right\}/\partial\varphi\left|{}_{\varphi=\hat{\varphi}^{0}_{\psi}}\right$.. We parameterize this line by $t\in\mathbb{R}$, namely $s(t)=s_{\psi}+t(s^{0}-s_{\psi})$. In particular, $s(0)=s_{\psi}$, corresponding to $H_{\psi}$, and $s(1)=s^{0}$, corresponding to the observed data. Then, the directional test for $s$ is conditional on being on the line $L^{*}_{\psi}$, i.e. conditional on s(t)/s(t)s(t)/$|$s(t)$|$. The directional $p$-value measuring the departure from $H_{\psi}$ along the line $L_{\psi}^{*}$ is defined as the probability that $s(t)$ is as far or farther from $s_{\psi}$ than is the observed value $s^{0}$:
@@
where the denominator is a normalizing constant. The upper limit $t_{\text{sup}}$ of the integrals in (7) is the largest value of $t$ for which the maximum likelihood estimate $\hat{\varphi}(t)$, corresponding to $s(t)$, exists. See sartori:2016 for more details.
In the particular case where $\psi$ and $\lambda$ are linear functions of the canonical parameter of an exponential family, the quantity (6) does not depend on $s$ and can therefore be ignored in computing the directional $p$-value (7). Moreover, the expected value $s_{\psi}$ simplifies to $s_{\psi}=\left\{-\ell_{\psi}^{0}\left(\hat{\varphi}^{0}_{\psi}\right)^{\top},{{0}_{q-d}^{\top}}\right\}^{\top}$.
If the original model $f(y;\theta)$ has not the exponential family form, then a tangent exponential model is used instead of (4). The construction of the tangent exponential model and its saddlepoint approximation are described in sartori:2016; see also davison2022tangent. Here this step is not needed since we work with the normal model which belongs to the exponential family, yet the saddlepoint approximation (5) will be used when $\psi$ is not a linear function of the canonical parameter, as in the heteroscedastic one-way MANOVA of Section 3.2.
@@ 2.3 Independent groups from multivariate normal distributions
Consider $g$ independent groups, and denote by $y_{ij}$ the independent observations from the $i$th group with multivariate normal distribution $N_{p}(\mu_{i},\Lambda_{i}^{-1})$, $i\in\{1,\dots,g\};j\in\{1,\dots,n_{i}\}$. The mean vectors $\mu_{i}$ and the concentration matrices $\Lambda_{i}$, symmetric and positive definite, are assumed unknown. Let $\text{tr}(A)$ indicate the trace of a square matrix $A$, and $\text{vec}(A)$ be the vector stacking the columns of $A$ one by one. We also define the vector $\text{vech}(A)$, obtained from $\text{vec}(A)$ by eliminating all upper triangular elements of $A$ when this is symmetric. These two vectors satisfy the relationship $D_{p}\text{vech}(A)=\text{vec}(A)$, where $D_{p}$ is the so-called duplication matrix (magnus1999, Section 3.8).
@@
where $\bar{y}_{i}=y_{i}^{\top}1_{n_{i}}/n_{i}$ with $1_{n_{i}}$ a $n_{i}$-dimensional vector of ones, $i\in\{1,\dots,g\}$. In this exponential family model, the canonical parameter is $\varphi=\{\xi_{1}^{\top},\dots,\xi_{g}^{\top},$ ${\text{vech}}(\Lambda_{1})^{\top},\dots,{\text{vech}}(\Lambda_{g})^{\top}\}^{\top}=\{\mu_{1}^{\top}\Lambda_{1},\dots,\mu_{g}^{\top}\Lambda_{g},$ ${\text{vech}}(\Lambda_{1})^{\top},\dots,{\text{vech}}(\Lambda_{g})^{\top}\}^{\top}$. The log-likelihood as a function of $\varphi$ is $\ell(\varphi)=\sum_{i=1}^{g}\ell_{i}(\varphi_{i})$ with $\varphi_{i}=\{\xi_{i}^{\top},{\text{vech}}(\Lambda_{i})^{\top}\}^{\top}$, and the $i$-th group’s contribution is
$\ell_{i}(\varphi_{i})$ $=n_{i}\xi_{i}^{\top}\bar{y}_{i}-\frac{1}{2}{\text{tr}}(\Lambda_{i}y_{i}^{\top}y_{i})+\frac{n_{i}}{2}\log|\Lambda_{i}|-\frac{n_{i}}{2}\xi_{i}^{\top}\Lambda_{i}^{-1}\xi_{i}$
$=\xi_{i}^{\top}n_{i}\bar{y}_{i}-\text{vech}(\Lambda_{i})^{\top}\left\{\frac{1}{2}D_{p}^{\top}D_{p}\text{vech}(y_{i}^{\top}y_{i})\right\}+\frac{n_{i}}{2}\log|\Lambda_{i}|-\frac{n_{i}}{2}\xi_{i}^{\top}\Lambda_{i}^{-1}\xi_{i}.$
@@
The score function is $\ell_{\varphi}(\varphi)=\left.{\partial}\ell(\varphi)\right/{\partial\varphi}=\{\ell_{\varphi_{1}}(\varphi_{1}),\dots,\ell_{\varphi_{g}}(\varphi_{g})\}^{\top}$, with
$\ell_{\varphi_{i}}(\varphi_{i})$ $=\left\{\ell_{\xi_{i}}(\varphi_{i})^{\top},\ell_{\text{vech}(\Lambda_{i})}(\varphi_{i})^{\top}\right\}^{\top}$
$=\left\{n_{i}\bar{y}_{i}^{\top}-n_{i}\xi_{i}^{\top}\Lambda_{i},\;\frac{n_{i}}{2}\text{vech}(\Lambda_{i}^{-1}-y_{i}^{\top}y_{i}/n_{i}+\Lambda_{i}^{-1}\xi_{i}\xi_{i}^{\top}\Lambda_{i}^{-1})^{\top}\right\}^{\top}.$
@@
The maximum likelihood estimates for $\mu_{i}$ and $\Lambda_{i}^{-1}$ are $\hat{\mu}_{i}=\bar{y}_{i}$ and $\hat{\Lambda}_{i}^{-1}=y_{i}^{\top}y_{i}/n_{i}-\bar{y}_{i}\bar{y}_{i}^{\top}$, respectively; hence, $\hat{\xi}_{i}=\hat{\Lambda}_{i}\hat{\mu}_{i}$, $i\in\{1,\dots,g\}$. Moreover, the observed information matrix $J_{\varphi\varphi}(\varphi)$ can be written in a block-diagonal form, with the group-specific diagonal block $J_{\varphi_{i}\varphi_{i}}(\varphi_{i})$, $i\in\{1,\dots,g\}$, equal to (huang2021directional, Section 2.3)
$J_{\varphi_{i}\varphi_{i}}(\varphi_{i})=\left[\begin{array}[]{cc}n_{i}\Lambda_{i}^{-1}&\quad-n_{i}(\xi_{i}^{T}\Lambda_{i}^{-1}\otimes\Lambda_{i}^{-1})D_{p}\\ -n_{i}D_{p}^{T}(\Lambda_{i}^{-1}\xi_{i}\otimes\Lambda_{i}^{-1})&\quad\frac{n}{2}D_{p}^{T}\{\Lambda_{i}^{-1}(I_{p}+2\xi_{i}\xi_{i}^{T}\Lambda_{i}^{-1})\otimes\Lambda_{i}^{-1}\}D_{p}\end{array}\right],$ (10)
where $\otimes$ denotes the Kronecker product (see, e.g., Lauritzen:1996, Section 5.1). Then, the determinant of $J_{\varphi\varphi}$ is such that $|J_{\varphi\varphi}(\varphi)|=\prod_{i=1}^{g}|J_{\varphi_{i}\varphi_{i}}(\varphi_{i})|\propto\prod_{i=1}^{g}|{\Lambda}_{i}^{-1}|^{p+2}$.